1.) A man is four times as old as his son. In three years,the father will be three times as old as his son. How old is each three years ago? Please show me the complete procedure,thank you.
Answer by mananth(16946)

son be x years
father = 4x

after 3 years
son will be x+3
father will be 4x+3
4x+3 = 3(x+3)
4x+3= 3x+9
4x-3x=9-3
x=6
son's ge = 6 now
father's age = 24 years
3 years ago
son = 3years
father = 21 years